Bob’s apple store utilizes a continuous review system. Current on had inventory is 295 bags of apples, with no open orders or backorders. He buys apples year round, 52 weeks a year, 5 days a week. He buys apples for $10.20 a bag. The following information is available about these bags of apples:

Demand = 80 bags/week

Order cost = $48/order

Annual Holding cost = 32 percent of the cost of a bag

Desired cycle-service level = 87 percent

Lead time = 2 weeks (10 working days)

Standard deviation of weekly demand = 12 bags

Go to two decimal points in all calculations. show work

3A. What is the annual holding cost (H) per unit expressed in dollars?                __________________ answer

(two decimal points)

3B. What is the annual demand (D) of bags of apples?                                        ___________________ answer

(two decimal points)

3C. What is order cost (S) expressed in dollars? (per order)                     ________----------------------------ans.

(two decimal points)

3D. What is the EOQ? ( I want a number here, not a noun)                                       ________________ answer           (two decimal points)                                  

3E. What would the average TBO (time between orders) (in weeks)?     _________________ answer (weeks)       (two decimal points)

3F. What is the safety stock? (I want a number here)                                        ___________________ answer

(two decimal points)

3G. What should R (re-order point) be?                 ________________ answer      (two decimal points)
Answer #1

3A.

Annual holding cost = 32%*purchase cost = 0.32*10.2 = $3.264 = $3.26 (Rounded to 2 decimal places)

3B

Annual Demand (D) = 80*52 = 4160

3C

Ordering cost per order = $48

3D

EOQ = Square root ((2*D*S)/h)   

EOQ = Square root ((2*4160*48)/3.26)

= Square-root(122503.0675)

=350.0043821

= 350 (Rounded to 2 decimal places)

3E

average TBO (time between orders) (in weeks) = No. of weeks / number of orders

NUmber of orders = Annual demand/EOQ = 4160/350 = 11.88571429

average TBO (time between orders) (in weeks) = 52/11.88571429 = 4.375 = 4.38 (Rounded to 2 decimal places)

3F

z value at 87% service level = 1.126391129

Standard deviation of demand during lead time = Standard deviation of weekly demand * Square root(lead time in weeks) = 12*Square root(2) = 16.97056275

Safety stock = z-value*Standard deviation of demand during lead time = 1.126391129*16.97056275 = 19.11549134 = 19.12 (Rounded to 2 decimal places)

3G

Re-order point = Demand per week * Lead time in week + Safety stock
= 80*2 + 19.12
= 179.12